From acd@roe.ac.uk Fri Feb 13 11:52:33 2015 Date: Thu, 12 Feb 2015 14:02:36 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tkerr@jach.hawaii.edu" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 30 January 2015 WFAU VDFS Science Archives progress meeting minutes, 30 January 2015 Present: NJC, RGM, ETWS, RPB, MAR, MSH, ACD Apologies: RSC, NCH, STV, KTN, AL DoNM: 10am Friday 13 February 2015 in the Villas meeting room, MAR in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: MAR to contact MJI about monthly WFAU/CASU telecons. Done. The telecons were now arranged and the first one had happened on Wednesday 28/1/15. Actions carried forward from 16/1/2015 meeting: ----------------------------------------------- ACTION: RPB (was MAR) to progress [was: 'have a look at'] the status of non-survey releases for registered projects. Continues. MAR had compiled a list of non-survey projects that could be released and sent it to RPB. RPB should progress these releases to completion. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this action still has low priority. ACTION: All to finish off VVV brainstorming tasks by end of January at the latest. Continues. NEW ACTION: ACD to schedule a meeting to review progress with the various tasks from the previous meetings, RSC to chair. ACTION: ETWS/NCH purchase a new license for mxODBC. Continues. Also the cost of this item should be established. ACTION ETWS/NJC: develop an automated tool to monitor progress in ingesting and curating surveys. Continues. ACTION: All to update the staff absences page. Continues. Many staff members had added their forthcoming absences but others had not. Specific points and new actions: -------------------------------- Project management: * Nothing to report. WFCAM, VISTA and VST updates: * Normal operations seemed to be continuing. Observing time on UKIRT was now divided amongst the various partners. Some time had been lost due to changing instruments on the telescope. Networking: * Nothing to report; everything appeared to be working satisfactorily. WSA/VSA/OSA/GES Operations: * ETWS reported: - The Chilean OSA data had been retransferred. These data are needed because they affect the zero points used to calibrate other surveys. ACTION: MAR check with Tom Shanks exactly what data are needed. - VSA/VVV P92 ingests were continuing. * RPB reported: - VMC P93 ingest had been started; currently the tiles were being created. - He was halfway through reading the UKIRT archive tapes (so far there was plenty of detail but not plot or character development). ACTION: ACD to check with JAC whether they have dispatched all the tapes or there are more to be sent. - VVV DR3 was still processing. Hardware and Systems: * MSH noted: - One NAS box memory had been failed and been replaced (2 TByte). Eclipse had sent an option to upgrade the capacity of the NAS boxes and a quote was expected next week. - We should have just about enough free disk space to last until April when it was anticipated money for new hardware would be available. Software: * NJC reported: - For OSA he had fixed problems related to the Chilean and UB filter data. - He had started working on the VMC and VVV ESO releases. ACTION: NJC noted that VISTA mosaicing crashes with normal filtering, which he should fix. * ETWS noted that he had fixed various bits and pieces, mostly concerned with both list-driven and non-list driven photometry. Review of Deadlines: * Progress towards forthcoming deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * The VIKING P92 Team release had been made ahead of schedule. ACTION: RPB to compile a Web page listing all the completed releases. Non-survey Data Release: * Nothing to report Astrogrid deployment, VO & Data Analysis services: * MSH reported that the UKIDSS DR10 GPS coverage maps had been published. Miscellaneous: * NJC noted that he had had a talk accepted for the forthcoming meeting on 'Milky Way Astrophysics from Wide-Field Surveys' to be held in the RAS premises in Burlington House on 30 March -1 April 2015. ACD, Original: 30/1/15, Revised: 30/1/15.